OpenStudy (anonymous):

Simplify the square root of (x+1)^4

OpenStudy (anonymous):

Given sqrt ( x + 1) = 4 We raise both sides to power 2 in order to clear the square root. [ sqrt ( x + 1) ] 2 = 4 2 and simplify x + 1 = 16 Solve for x. x = 15 NOTE: Since we squared both sides without putting any conditions, extraneous solutions may be introduced, checking the solutions is necessary. Left side (LS) of the given equation when x = 15 LS = sqrt (x + 1) = sqrt (15 + 1) = 4 Right Side (RS) of the given equation when x = 15 RS = 4 For x = 15, the left and the rigth sides of the given equation are equal: x = 15 is a solution to the given equation. Example 2 : Find all real solutions to the equation

OpenStudy (anonymous):

(x+1) is raised to four

OpenStudy (anonymous):

what was all that gibberish?

OpenStudy (anonymous):

the square root of something raised to the power of four is that something raised to the power of 2

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

can you make it clear?

OpenStudy (anonymous):

lol it broke it all down

OpenStudy (anonymous):

wait this is the given \[\sqrt{(x+1)^4} \]

OpenStudy (anonymous):

yeah my bad

OpenStudy (anonymous):

@No.name your real funny im trying to help her

OpenStudy (anonymous):

\[\huge [(x+1)^2]^2\] right?

OpenStudy (anonymous):

stop

OpenStudy (anonymous):

\[\sqrt{x^4}=x^2\\ \sqrt{(x+1)^4}=(x+1)^2\]

OpenStudy (anonymous):

so i was right

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

lol